Given: K(0, 4), A(3, 6), T(6, 2), E(0, -2)
Prove: KATE is a trapezoid
Find the slope of all 4 sides- show all work

Answer by Alan3354(69443)

Given: K(0, 4), A(3, 6), T(6, 2), E(0, -2)
Prove: KATE is a trapezoid
----------------
Find the slope of KA
Slope m = diffy/diffx = (6-4)/(3-0) = 2/3
-------
Do the other 3 sides the same way.
If 2 sides have the same slope, they're parallel --> a trapezoid.